# Distributed Computing From First Principles

### v1

@@ 7.2 Code-Level Verification
Code-level verification takes a different approach compared to traditional model-level formal verification, where an abstract model of the problem space is created and checked for violations of safety and liveness properties. Instead, code-level verification directly examines the source code of the system, often involving running MPI/C source code within a specialized verification scheduler. A significant challenge in verifying concurrent programs is the potentially exponential number of possible interleavings of the execution of parallel threads or processes. To address this, code-level verification techniques often employ partial order reduction (POR) algorithms to identify and eliminate redundant or equivalent interleavings, thereby significantly reducing the state space that needs to be explored [46].
@@ 8.2 Storage Systems
Every volume, despite its sophisticated architecture, includes the following core features: a data file (the actual stored data), an index file (a searchable data structure), a journal file (which allows for persistence between restarts), a checkpoint (a snapshot that provides safe points for recovery after failures), and a client (an application that has access to the file system). These features have various names across different file systems, such as HDFS [40], Google File System [26], and Tectonic [36]. Some file systems [38] allow for sequential modification to a disk in a log-like manner. Another way to categorize storage systems is as single-tenant (Haystack [15], F4 [35], HDFS [40]) or multi-tenant (Tectonic [36]). A number of these storage systems utilize variants of Paxos and Raft to achieve distributed consensus across volumes. For more information on distributed consensus algorithms, please refer to Chapter 4, Sections 4.1 and 4.3.
@@
There is room for optimizations to improve storage, leading to the development of some distributed storage systems. Haystack [15] improves upon the Unix file system by saving metadata in memory, thereby making lookups faster, while reading the actual file saved in the volume involves disk operations, thereby increasing throughput. The metadata is spread over a wide area of memory, allowing for compact representation in the Haystack store, rather than each directory storing the metadata. However, it is necessary to ensure that there is sufficient memory to load a significant amount of the metadata for lookup purposes. When an object is saved in a volume in a Haystack store, it is replicated to many secondary volumes for redundancy. The Haystack directory maps each volume to the metadata to facilitate reading the actual file using the metadata. The Haystack cache is used to increase the lookup speed of popular objects. This is ideal for immutable blobs, hence it follows the philosophy of write once, read always, and rarely update.
@@
Haystack is deficient in utilizing the access patterns of requests for objects, and even the Haystack cache is not sufficient for granular control of access patterns. Hence, F4 [35] was developed to allow for incorporating access patterns into the design of a store. Rather than a RAID-like replication scheme, F4 utilizes distributed erasure coding and Reed-Solomon codes to ensure smarter, lower-cost replication to several secondary volumes. Objects saved in the store begin as ”hot” and then ”warm” over time. The idea is based on the pattern that newer objects are more likely to be accessed, while older objects are less likely to be accessed. This concept is known as temperature zones. Volumes are structured in a way that once a memory limit for a volume is reached, it is closed for writing and becomes read-only. These volumes are organized into temperature zones to take advantage of access patterns. There are also many systems (Pelican [14], RADOS [48]) that allow for intelligent use of access patterns, and a system (CRUSH [47]) that provides the option to specify weights for access patterns. Tectonic [36] is an evolution of the Haystack storage system, as it introduced multi-tenant features. Each tenant has its own namespace and delegates metadata lookup to a dedicated key store.
@@ 8.6 Exercises for the Readers
    As a helpful starting point, consider the strategies for parallelizing linear algebra computations, as presented in Jeff Dean’s bachelor thesis available at Jeff Dean’s Bachelor thesis. These techniques for distributing and combining computations can offer valuable insights 111 For potential future development and expansion of this project, consider exploring the following ideas: including asynchronous late update and dropout 222 ©Kenneth Emeka Odoh [Uncaptioned image]. All contents of this book is covered under [Uncaptioned image] license

### v2

@@ 7.2 Code-Level Verification
Code-level verification takes a different approach compared to traditional model-level formal verification, where an abstract model of the problem space is created and checked for violations of safety and liveness properties. Instead, code-level verification directly examines the source code of the system, often involving running MPI/C source code within a specialized verification scheduler. A significant challenge in verifying concurrent programs is the potentially exponential number of possible interleavings of the execution of parallel threads or processes. To address this, code-level verification techniques often employ partial order reduction (POR) algorithms to identify and eliminate redundant or equivalent interleavings, thereby significantly reducing the state space that needs to be explored [Vo2009].
@@ 8.2 Storage Systems
Every volume, despite its sophisticated architecture, includes the following core features: a data file (the actual stored data), an index file (a searchable data structure), a journal file (which allows for persistence between restarts), a checkpoint (a snapshot that provides safe points for recovery after failures), and a client (an application that has access to the file system). These features have various names across different file systems, such as HDFS [shvachko2010hadoop], Google File System [ghemawat2003google], and Tectonic [PanTectonic2021]. Some file systems [rosenblum1992] allow for sequential modification to a disk in a log-like manner. Another way to categorize storage systems is as single-tenant (Haystack [Beaver2010], F4 [Muralidharf42014], HDFS [shvachko2010hadoop]) or multi-tenant (Tectonic [PanTectonic2021]). A number of these storage systems utilize variants of Paxos and Raft to achieve distributed consensus across volumes. For more information on distributed consensus algorithms, please refer to Chapter 4, Sections 4.1 and 4.3.
@@
There is room for optimizations to improve storage, leading to the development of some distributed storage systems. Haystack [Beaver2010] improves upon the Unix file system by saving metadata in memory, thereby making lookups faster, while reading the actual file saved in the volume involves disk operations, thereby increasing throughput. The metadata is spread over a wide area of memory, allowing for compact representation in the Haystack store, rather than each directory storing the metadata. However, it is necessary to ensure that there is sufficient memory to load a significant amount of the metadata for lookup purposes. When an object is saved in a volume in a Haystack store, it is replicated to many secondary volumes for redundancy. The Haystack directory maps each volume to the metadata to facilitate reading the actual file using the metadata. The Haystack cache is used to increase the lookup speed of popular objects. This is ideal for immutable blobs, hence it follows the philosophy of write once, read always, and rarely update.
@@
Haystack is deficient in utilizing the access patterns of requests for objects, and even the Haystack cache is not sufficient for granular control of access patterns. Hence, F4 [Muralidharf42014] was developed to allow for incorporating access patterns into the design of a store. Rather than a RAID-like replication scheme, F4 utilizes distributed erasure coding and Reed-Solomon codes to ensure smarter, lower-cost replication to several secondary volumes. Objects saved in the store begin as ”hot” and then ”warm” over time. The idea is based on the pattern that newer objects are more likely to be accessed, while older objects are less likely to be accessed. This concept is known as temperature zones. Volumes are structured in a way that once a memory limit for a volume is reached, it is closed for writing and becomes read-only. These volumes are organized into temperature zones to take advantage of access patterns. There are also many systems (Pelican [Balakrishnan2014], RADOS [weil2007]) that allow for intelligent use of access patterns, and a system (CRUSH [weil2006]) that provides the option to specify weights for access patterns. Tectonic [PanTectonic2021] is an evolution of the Haystack storage system, as it introduced multi-tenant features. Each tenant has its own namespace and delegates metadata lookup to a dedicated key store.
@@ 8.6 Exercises for the Readers
    As a helpful starting point, consider the strategies for parallelizing linear algebra computations, as presented in Jeff Dean’s bachelor thesis available at Jeff Dean’s Bachelor thesis. These techniques for distributing and combining computations can offer valuable insights  [^1] For potential future development and expansion of this project, consider exploring the following ideas: including asynchronous late update and dropout 222©Kenneth Emeka Odoh [Uncaptioned image]. All contents of this book is covered under [Uncaptioned image] license